# Spiking GS: Towards High-Accuracy and Low-Cost Surface Reconstruction via Spiking Neuron-based Gaussian Splatting

### v3

@@ Abstract
3D Gaussian Splatting is capable of reconstructing 3D scenes in minutes. Despite recent advances in improving surface reconstruction accuracy, the reconstructed results still exhibit bias and suffer from inefficiency in storage and training. This paper provides a different observation on the cause of the inefficiency and the reconstruction bias, which is attributed to the integration of the low-opacity parts (LOPs) of the generated Gaussians. We show that LOPs consist of Gaussians with overall low-opacity (LOGs) and the low-opacity tails (LOTs) of Gaussians. We propose Spiking GS to reduce such two types of LOPs by integrating spiking neurons into the Gaussian Splatting pipeline. Specifically, we introduce global and local full-precision integrate-and-fire spiking neurons to the opacity and representation function of flattened 3D Gaussians, respectively. Furthermore, we enhance the density control strategy with spiking neurons’ thresholds and a new criterion on the scale of Gaussians. Our method can represent more accurate reconstructed surfaces at a lower cost. The supplementary material and code are available at https://github.com/zju-bmi-lab/SpikingGS.
@@
1 Introduction
3D surface reconstruction from multiview RGB images is a challenging task in computer graphics and vision Turkulainen et al. (2024); Verbin et al. (2022). Methods based on Neural Radiance Fields (NeRF) Mildenhall et al. (2021); Guo et al. (2022); Martin-Brualla et al. (2021); Niemeyer et al. (2022); Wang et al. (2021) can extract geometry information from an implicit representation but require extensive computational time for training, limiting their application scenarios. In contrast, 3D Gaussian Splatting (3DGS) Kerbl et al. (2023) is exceptionally fast in 3D reconstruction, which is gaining more and more attention.
[…]
In this paper, we show a detailed analysis (Sec. 3.2) about above observations in Fig. 1 (a) and Fig. 1 (c), based on which we further identify two primary sources of LOPs: low-opacity Gaussians (LOGs) and the low-opacity tails (LOTs) of Gaussians (Fig. 1 (b)). To handle LOGs and LOTs, we introduce a global full-precision integrate-and-fire (FIF) spiking neuron on the opacities of Gaussians and a local FIF spiking neuron considering the Gaussian representation itself, respectively. The former uses a learnable threshold to raise the lower bound of the opacity value of the scene, which effectively reduces the number of LOGs; the latter introduces the discontinuity in the Gaussian representation function, which reduces the contribution of LOTs in 3DGS integration. By suppressing the integration of LOPs for view rendering, we can adaptively regulate LOPs, facilitating robust convergence towards the optimal scenario for various scenes. Although Spiking GS excels in surface reconstruction, it may overly remove Gaussians in occluded or sparsely observed regions, which slightly harms the reconstruction accuracy. To avoid this, we introduce the scale-based clone to compensate for the missing Gaussians in these regions. Moreover, we implement several regularization losses to further enhance method’s performance. Extensive experiments show that Spiking GS achieves comparable or superior 3D reconstruction performance on mainstream datasets with much fewer Gaussians and significantly outperforms other methods in semi-transparent scenarios. Note that Spiking GS does not rely on any priors from pre-trained geometry (depth or normal) estimation models. A glimpse of Spiking GS’s performance is shown in Fig. 1 (d). In summary, we make the following contributions:
  • We introduce a global FIF spiking neuron on the opacity of Gaussians and a local FIF spiking neuron on each Gaussian representation function, which reduces the number of LOPs and suppresses the integration of LOPs for view rendering.
  • We propose optimized training strategies, including the scale-based clone and several regularization losses to improve reconstruction accuracy. Without priors from pre-trained geometry estimation models, the proposed Spiking GS achieves higher reconstruction accuracy with fewer Gaussians for low storage occupation and high optimization efficiency.
2 Related Work
Surface Reconstruction with implicit representation. Neural Radiance Field (NeRF, Mildenhall et al. 2021) has explored a novel way to reconstruct 3D surface based on a neural radiance field. However, such reconstruction suffers from significant artifacts due to a lack of constraints on geometry Verbin et al. (2022); Wang et al. (2021); Yariv et al. (2021); Fu et al. (2022); Wang et al. (2022; 2023) and defects of MLP in representing real-world scene Liao et al. (2024). To improve the reconstruction quality, some of the subsequent works focus on modifying the MLP-based representation by either integrating with Signed Distance Functions (SDFs), a more explicit surface representation Wang et al. (2021); Yariv et al. (2021); Fu et al. (2022); Wang et al. (2022; 2023), or introducing discontinuities into the density field Liao et al. (2024). Other works adopt strong priors to regularize the reconstruction, such as priors from depth Deng et al. (2022); Yu et al. (2022) and MVS Zhang et al. (2021); Zuo & Deng (2022). Nevertheless, these methods still require prolonged time for optimizing the neural radiance fields. In contrast, our method is developed on the fast 3DGS method to achieve more accurate surface reconstruction with lower cost.
Surface Reconstruction with 3D Gaussians. Although 3D Gaussian Splatting Kerbl et al. (2023) achieves high fidelity on novel view synthesis, it struggles to reconstruct accurate surface due to the difficulty of aligning primitives with the actual surface Dai et al. (2024) and the inconsistency between mesh extraction and view rendering Yu et al. (2024). To address the issues,  Chen et al. 2024b; Guédon & Lepetit 2024 utilize a regularization to flatten the 3D Gaussian ellipsoids into flat shapes, while Dai et al. 2024; Huang et al. 2024 directly use flattened 3D Gaussians as primitives. Others redesigned the opacity modeling of Gaussians Lyu et al. (2024); Yu et al. (2024), improving the details of the reconstructed surfaces. However, these approaches still face challenges in the accuracy of surface reconstruction in scenes including objects with complex materials (i.e, semi-transparent). The proposed Spiking GS identifies a widely overlooked correlation between the reconstruction bias and LOPs, achieving considerable improvement on the reconstruction quality and efficiency by reducing the number of LOPs.
Efficient Geometry Representation of 3D Gaussians. The use of a large number of Gaussians to accurately represent scenes often results in inefficient training and storage. To address this, recent efforts have focused on utilizing compact feature encodings and pruning strategies to improve efficiency in Gaussian representations Chen et al. (2024c); Lee et al. (2024); Lu et al. (2024). However, these methods sacrifice surface reconstruction accuracy. Other works Dai et al. (2024); Guédon & Lepetit (2024) eliminate the occurrence of LOGs or introduce opacity-based regularization terms to reduce the extra Gaussians. However, they suffer from limited representation capabilities, particularly in scenes with semi-transparent objects, or limited effect on the occurrence of Gaussians. Although our method shares a similar target of reducing the number of Gaussians for improved efficiency, we specifically focus on minimizing the number of LOPs in integration. As a result, we not only effectively achieve the target but also maintain the representation capabilities for accurate surface reconstruction in various scenes.
Spiking Neural in Surface Reconstruction. Spiking Neural Networks (SNNs) have received extensive attention recently in many fields due to their energy efficiency Kim et al. (2022); Ren et al. (2024). It has been utilized in processing video or event signals Paredes-Vallés & De Croon (2021); Zhu et al. (2022), classification tasks Li et al. (2022); Zhang et al. (2022), and adversarial attacks Ding et al. (2022); Kundu et al. (2021); Liang et al. (2022); Zhou et al. (2021). Recently, Liao et al. (2024) introduce the spiking neuron into the NeRF-based neural surface reconstruction for better performance in representing discontinuities of the scene. The proposed spiking GS is also equipped with the spiking neurons but for its efficiency in removing the LOPs in 3DGS. Experiments show that it significantly improves the performance of 3DGS-based surface reconstruction.
3 Preliminaries and Analysis
In this section, we present preliminaries in Spiking GS, including flattened Gaussian splatting and spiking neurons. We then conduct an analysis of LOPs, which is the core of Spiking GS.
3.1 Preliminary
Flattened Gaussian Splatting. 3DGS Kerbl et al. (2023) use a set of 3D Gaussian primitives $\mathcal{G}$ to model the real world scene, each of which is parameterized by an opacity $\alpha\in[0,1]$, central point’s position $\mathbf{p}\in\mathbb{R}^{3\times 1}$, scale vector $\mathbf{S}\in\mathbb{R}^{3\times 1}$, and the rotation matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$ parameterized by a quaternion $r$. In this paper, we apply flattened 3D Gaussians Dai et al. (2024) as the 3D primitive due to its ability to reconstruct richer surface details than other methods. The distribution of flattened 3D Gaussian is defined by:
where $\mathrm{x}$ is the 3D coordinate, $\Sigma$ is the covariance matrix expressed as $\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}$, $\mathbf{S}$ is set as $[\mathbf{S}^{\mathrm{x}},\mathbf{S}^{\mathrm{y}},0]^{\top}$ Dai et al. (2024). In the rendering process, flattened 3D Gaussians in the world coordinates are transformed into the camera coordinates using a world-to-camera transformation matrix $\mathbf{W}$ and projected onto the image plane using an affine approximation matrix $\mathbf{J}$ Zwicker et al. (2001):
where the third row of $\Sigma$ is skipped for projection. After that, the rendered color of each pixel is calculated via alpha-blending of $N$ ordered Gaussians overlapped on pixel points $\mathbf{u}$:
where $\mathbf{c}(\mathbf{u})$ is the pixel value at point $\mathbf{u}$, $\mathbf{c}_{i}$ is the RGB color of Gaussians, $\mathcal{G}_{i}^{\prime}\left(\mathbf{u}\right)=e^{-\frac{1}{2}(\mathbf{u}-\mathbf{p})^{\top}\Sigma^{\prime-1}(\mathbf{u}-\mathbf{p})}$ is the Gaussians projected in 2D, $\alpha_{i}$ is the Gaussian opacity, $\omega_{i}$ is the weighted Gaussian opacity. During training, Gaussians’ attributes are optimized by a photometric loss, and the number of Gaussians is controlled by an adaptive density control process Kerbl et al. (2023).
@@ 3.1 Preliminary
Spiking Neuron. We introduce the simplified full-precision integrate-and-fire (FIF) spiking neuron Li et al. (2022); Liao et al. (2024). The model is described as:
$O=I\cdot s,~{}~{}~{}s=\begin{cases}0&I<\bar{V},\\ 1&\text{otherwise},\end{cases}$ (4)
@@
where $I$ is the input value, $\bar{V}$ is a spiking threshold, $O$ is the output value. To make the spiking neuron differentiable, we use the surrogate gradient Li et al. (2021); Liao et al. (2024) to calculate the gradient during training as:
@@ 3.2 Analyzing Low-opacity Parts
The suboptimal scenario leads to the emergence of excessive LOPs. As shown in Eq. (3), the pixel value in a given view is computed by integrating the color and opacity of the nearby projected Gaussians. If we focus on reducing the loss between the rendered and ground truth pixel values $\mathbf{c}(\mathbf{u})$ while overlooking regularizing the contribution of each Gaussian, there exists two possible scenarios for deriving the same pixel value $\mathbf{c}(\mathbf{u})$: the scenario where only a few high-opacity Gaussians are involved, or the scenario where both a few high-opacity Gaussians and excessive low-opacity Gaussians are involved. According to Dai et al. (2024); Guédon & Lepetit (2024), the latter is suboptimal. We observe that existing state-of-the-art (SOTA) methods tend to converge to such suboptimal scenario. We validate this by visualizing the representative pattern of the $\omega$ values along the view rays generated by SOTA methods in Fig. 2 (a) and (b), showing that excessive LOPs are introduced by excessively generated Gaussians in the SOTA methods. This is further validated by the statistic results in Fig. 2 (c) and (d).
Excessive LOPs bring the bias of surface reconstruction as extracting surfaces from Gaussians involves similar opacity integration in Eq. (3). Take the widely used TSDF fusion Huang et al. (2024) as an example. It first computes the depth for every pixel point $\mathbf{u}$ as:
@@
where $t_{i}$ is the depth of the $i$-th 3D Gaussians. According to Eq. (6), the computed depth may be biased by the redundant LOPs. We visualize the estimated bias in Fig. 2 (a) and (b). The error in the depth map will be passed to the extracted mesh since the TSDF fusion method fuses depth maps from different views to generate the final mesh. Such an issue also exists in other advanced surface extraction methods, e.g., the tetrahedral-based surface extraction proposed in Yu et al. (2024), due to their similar computation method in Eq. (6) for the opacity field derivation. Although we have shown that numerous LOPs bring the bias of surface reconstruction, we would like to emphasize that they are not the only cause, i.e, fewer LOPs in the integration is one of the necessary conditions towards the optimal Gaussians for 3DGS-based methods. Due to a training process driven by photometric loss instead of direct regularization on geometry, it is possible that the reconstructed surfaces still has a high reconstruction bias even with fewer LOPs. For example, G-Surfel sometimes shows a high depth error but contains fewer LOPs compared to other methods (Fig. 2, (a)).
@@
Figure 2: Visualization of $\omega$ value on a view ray of (a) Drums and (b) Mic from NeRF-Synthetic Dataset Mildenhall et al. (2021). We select representative patterns of $\omega$ near the actual surface. The height of each bin represents the $\omega$ value, the red and green lines represent the GT depth and the estimated depth, respectively, which are the same as Fig. 1 (a). We compare five methods (3DGS, 2DGS, G-Surfel, GOF, and ours) and the optimal case. Short bins indicate the presence of LOPs spreading out around the actual surface positions. We also mark the corresponding depth errors (the absolute difference between the ground truth depth and predicted depth) behind the name of each method. Statistics on two types of LOPs on (c) Drums, (d) Mic from NeRF-Synthetic Dataset, including the number of Gaussians (total height of the colored bar), the proportion of LOGs (light orange section), and the number of LOTs (blue bar, count as the averaged overlaps of Gaussian’s tails with a Gaussian representation function value below 0.1 in multiple views).

### v5

@@ Abstract
3D Gaussian Splatting is capable of reconstructing 3D scenes in minutes. Despite recent advances in improving surface reconstruction accuracy, the reconstructed results still exhibit bias and suffer from inefficiency in storage and training. This paper provides a different observation on the cause of the inefficiency and the reconstruction bias, which is attributed to the integration of the low-opacity parts (LOPs) of 3D Gaussians. We identify the source of LOPs and propose Spiking GS to reduce two types of LOPs by integrating spiking neurons into the Gaussian Splatting pipeline. Specifically, we introduce global and local full-precision integrate-and-fire spiking neurons to the opacity and representation function of flattened 3D Gaussians, respectively. Furthermore, we enhance the density control strategy with spiking neurons’ thresholds and a new cloning criterion. Our method can represent more accurate reconstructed surfaces at a lower cost. The code is available at https://github.com/zju-bmi-lab/SpikingGS.
Figure 1: Left: we show a visual comparison of the extracted mesh for the Mic object from the NeRF-Synthetic dataset [29]. We additionally compare the Gaussians’ number (#G), number of Low-Opacity Gaussians (#LOGs), Chamfer Distance (CD), and optimization time between our and previous methods. We regard opacity lower than 0.2 as low-opacity for statistics. Right: we show statics of different methods on the NeRF-Synthetic dataset [29] in terms of CD, #G, and training time visualized by the size of the circles, bigger circles indicate faster speed; numbers indicate the specific time cost in minutes.
1 Introduction
3D surface reconstruction from multiview RGB images is a challenging task in computer graphics and vision [34, 35]. Methods based on Neural Radiance Fields (NeRF) [29, 13, 28, 30, 36] can extract geometry information from an implicit representation but require extensive computational time for training, limiting their application scenarios. In contrast, 3D Gaussian Splatting (3DGS) [18] is exceptionally fast in 3D reconstruction, which is gaining more and more attention.
In particular, 3DGS-based methods integrate the opacity of Gaussians for view rendering and surface reconstruction. In the optimal scenario [5, 12], only a few high-opacity Gaussian tightly clustered around the surfaces contribute to integration. However, we observe that a suboptimal situation often arises, where Gaussians with low- and high-opacity are distributed around the surfaces, yielding the emergence of excessive Low-Opacity Parts (LOPs) in the 3DGS integration. Failing to suppress the emergence of LOPs results in excessive 3D Gaussians, hindering training and storage efficiency. Besides, LOPs also negatively influence surface reconstruction, leading to surface bias. Most of the previous 3DGS-based surface reconstruction methods overlook the effects of LOPs, focusing instead on improving the reconstructed surface’s quality through extra regularization [5], enhancing mesh extraction method [41], and new Gaussian primitives [2, 5, 14]. Despite the effectiveness of these methods, there is still considerable potential to improve the accuracy and efficiency of reconstruction in the aspect of suppressing excessive LOPs.
In this paper, we conduct a detailed analysis of the above observations in Sec. 3.2, based on which we further identify two primary sources of LOPs: Low-Opacity Gaussians (LOGs) and the Low-Opacity Tails (LOTs) of Gaussians (intuitive interpretation shown in Fig. 2). To handle LOGs and LOTs, we implement a global full-precision integrate-and-fire (FIF) spiking neuron on the opacities of Gaussians and a local FIF spiking neuron considering the Gaussian representation itself, respectively. The former uses a learnable threshold to raise the lower bound of the opacity value of the scene, which effectively reduces the number of LOGs; the latter introduces the discontinuity in the Gaussian representation function, which reduces the contribution of LOTs in 3DGS integration. By suppressing the integration of LOPs for view rendering, we can adaptively regulate LOPs, facilitating robust convergence towards the optimal scenario for various scenes. Although Spiking GS excels in surface reconstruction, it may overly remove Gaussians in occluded regions, which slightly harms the reconstruction accuracy. To avoid this, we introduce the scale-based clone to compensate for the missing Gaussians in these regions. Moreover, we implement several regularization losses to further enhance the method’s performance. Extensive experiments show that Spiking GS achieves comparable or superior 3D reconstruction performance on mainstream datasets with much fewer Gaussians and significantly outperforms other methods in semi-transparent scenarios. A glimpse of Spiking GS’s performance is shown in Fig. 1. In summary, we make the following contributions:
  • We implement a global FIF spiking neuron on the opacity of Gaussians and a local FIF spiking neuron on each Gaussian representation function, which reduces LOPs.
  • We propose a scale-based clone method and several regularization losses to improve reconstruction accuracy. The proposed Spiking GS achieves higher reconstruction accuracy with fewer Gaussians for low storage occupation and high optimization efficiency.
2 Related Work
Surface Reconstruction with Implicit Representation. Neural Radiance Field [29] has explored a novel way to reconstruct 3D surface based on a neural radiance field. However, such reconstruction suffers from significant artifacts due to insufficient geometric constraints [35, 36, 39, 11, 37, 38] and defects of MLP in representing real-world scene [25, 32]. To improve the reconstruction quality, some of the subsequent works focus on modifying the MLP-based representation by either integrating with Signed Distance Functions (SDFs), a more explicit surface representation [36, 39, 11, 37, 38], or introducing discontinuities into the density field [25, 32]. Other works adopt strong priors to regularize the reconstruction, such as priors from depth [6, 40] and MVS [42, 46]. Nevertheless, these methods still require prolonged time for optimizing the neural radiance fields. In contrast, our method is developed on the fast 3DGS method to achieve more accurate surface reconstruction with lower cost.
Surface Reconstruction with 3D Gaussians. Although 3D Gaussian Splatting [18] achieves high fidelity on novel view synthesis, it struggles to reconstruct accurate surface due to the difficulty of aligning primitives with the actual surface [5] and the inconsistency between mesh extraction and view rendering [41]. To address these issues,  [3, 12] utilize a regularization to flatten the 3D Gaussian ellipsoids into flat shapes, while [5, 14] directly use flattened 3D Gaussians as primitives. Others redesigned the opacity modeling of Gaussians [27, 41], improving the details of the reconstructed surfaces. However, these approaches still face challenges in the accuracy of surface reconstruction in scenes including objects with complex materials (i.e, semi-transparent). The proposed Spiking GS identifies a widely overlooked correlation between the reconstruction bias and LOPs, achieving considerable improvement on the reconstruction quality and efficiency by reducing LOPs.
Efficient Geometry Representation of 3D Gaussians. The use of a large number of Gaussians to accurately represent scenes often results in inefficient training and storage. To address this, recent efforts have focused on utilizing compact feature encodings and pruning strategies to improve efficiency in Gaussian representations [4, 21, 26]. However, these methods compromise surface reconstruction accuracy. Other works [5, 12] eliminate the occurrence of LOGs or introduce opacity-based regularization terms to reduce the extra Gaussians. However, they suffer from limited representation capabilities, particularly in scenes with semi-transparent objects, or limited effect on the occurrence of Gaussians. Although our method shares a similar target of reducing the number of Gaussians for improved efficiency, we specifically focus on minimizing the number of LOPs in integration. As a result, we not only effectively achieve the target but also maintain the representation capabilities for accurate surface reconstruction in various scenes.
Spiking Neurons in Surface Reconstruction. Spiking Neural Networks (SNNs) have received extensive attention recently in many fields due to their energy efficiency [19, 33]. It has been utilized in processing video or event signals [31, 45], classification tasks [22, 43], and adversarial attacks [7, 20, 24, 44]. Recently, Liao et al. [25] introduce the spiking neuron into the NeRF-based neural surface reconstruction for better performance in representing discontinuities of the scene. The proposed spiking GS is also equipped with the spiking neurons but for its efficiency in removing the LOPs in 3DGS. Experiments show that it significantly improves the performance of 3DGS-based surface reconstruction.
Figure 2: Illustration of LOPs (LOGs and LOTs). From top to bottom, the first (second) row shows the front (side) view of the Gaussians around the surface.
3 Preliminaries and Analysis
In this section, we present preliminaries in Spiking GS, including flattened Gaussian splatting and spiking neurons. We then conduct an analysis of LOPs.
3.1 Preliminary
Flattened Gaussian Splatting. 3DGS [18] use a set of 3D Gaussian primitives $\mathcal{G}$ to model the real world scene, each of which is parameterized by an opacity $\alpha\in[0,1]$, central point’s position $\mathbf{p}\in\mathbb{R}^{3\times 1}$, scale vector $\mathbf{S}\in\mathbb{R}^{3\times 1}$, and the rotation matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$ parameterized by a quaternion $\mathbf{r}$. In this paper, we apply flattened 3D Gaussians [5] as the 3D primitive due to its ability to reconstruct richer surface details than other methods. The distribution of flattened 3D Gaussian is defined by:
where $\mathbf{x}$ is the 3D coordinate, $\Sigma$ is the covariance matrix expressed as $\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}$, set as $[\mathbf{S}^{\mathrm{x}},\mathbf{S}^{\mathrm{y}},0]^{\top}$ [5]. In the rendering process, Gaussians in the world coordinates are transformed into the camera coordinates using a world-to-camera transformation matrix $\mathbf{W}$ and projected onto the image plane using an affine approximation matrix $\mathbf{J}$ [47]:
where the third row and column of $\Sigma$ is skipped for projection. After that, the rendered color of each pixel is calculated via alpha-blending of $N$ ordered Gaussians overlapped on pixel points $\mathbf{u}$:
where $\mathbf{c}(\mathbf{u})$ is the pixel value at point $\mathbf{u}$, $\mathbf{c}_{i}$ is the RGB color of Gaussians, $\mathcal{G}_{i}^{\prime}\left(\mathbf{u}\right)=e^{-\frac{1}{2}(\mathbf{u}-\mathbf{p})^{\top}\Sigma^{\prime-1}(\mathbf{u}-\mathbf{p})}$ is the Gaussians projected in 2D, $\alpha_{i}$ is the Gaussian opacity, $\omega_{i}$ is the weighted Gaussian opacity. During training, Gaussians’ attributes are optimized by a photometric loss, and the number of Gaussians is controlled by an adaptive density control process [18].
@@ 3.1 Preliminary
Figure 4: Visualization of $\omega$ value on a view ray of (a) Drums and (b) Mic from NeRF-Synthetic Dataset [29]. We select representative patterns of $\omega$ near the actual surface. The height of each bin represents the $\omega$ value, the red and green lines represent the GT depth and the estimated depth. Short bins indicate the presence of LOPs spreading out around the actual surface positions. Numbers indicate the corresponding depth errors. Statistics on two types of LOPs on (c) Drums, (d) Mic from NeRF-Synthetic Dataset, including the number of Gaussians, the proportion of LOGs, and the number of LOTs (averaged overlaps of Gaussian’s tails with a Gaussian representation function value below 0.1 in multiple views).
@@
Spiking Neuron. We implement a simplified full-precision [^1] Following previous work, we note the spiking neuron with float rather than binary output as “full-precision” spiking neuron. integrate-and-fire (FIF) spiking neuron [22, 25]. The model is described as:
$o=ls(l),~{}~{}~{}s=\begin{cases}0&l<\bar{V},\\ 1&\text{otherwise},\end{cases}$ (4)
@@
where $l$ is the input value, $\bar{V}$ is a spiking threshold, $o$ is the output value. To make the spiking neuron differentiable, we use the surrogate gradient [23, 25] to calculate the gradient during training as:
@@ 3.2 Analyzing Low-opacity Parts
The suboptimal scenario leads to the emergence of excessive LOPs. As shown in Eq. (3), the pixel value in a given view is computed by integrating the color and opacity of the nearby projected Gaussians. If we focus on reducing the loss between the rendered and ground truth pixel values $\mathbf{c}(\mathbf{u})$ while overlooking regularizing the contribution of each Gaussian, there exist two possible scenarios for deriving the same pixel value $\mathbf{c}(\mathbf{u})$: the scenario where only a few high-opacity Gaussians are involved, or the scenario where both a few high-opacity Gaussians and excessive low-opacity Gaussians are involved. According to [5, 12], the latter is suboptimal. We observe that existing state-of-the-art (SOTA) methods tend to converge to such suboptimal scenarios, as illustrated in Fig. 3. We validate this observation by visualizing the representative pattern of the $\omega$ values along the view rays generated by SOTA methods in Fig. 4 (a) and (b), showing that excessive LOPs are introduced in the SOTA methods. This is further validated by the statistic results in Fig. 4 (c) and (d).
Excessive LOPs bring the bias of surface reconstruction as extracting surfaces from Gaussians involves similar opacity integration in Eq. (3), as shown in Fig. 3. Take the widely used TSDF fusion [14] as an example. It first computes the depth for every pixel point $\mathbf{u}$ as:
@@
where $t_{i}$ is the depth of the $i$-th 3D Gaussians. According to Eq. (6), the computed depth may be biased by redundant LOPs. We visualize the estimated bias along the ray in Fig. 4 (a) and (b). The error in the depth map will be passed to the extracted mesh since the TSDF fusion method fuses depth maps from different views to generate the final mesh. Such an issue also exists in other advanced surface extraction methods, e.g., the tetrahedral-based surface extraction proposed in [41], due to their similar computation method in Eq. (6) for the opacity field derivation. Although we have shown that numerous LOPs bring the bias of surface reconstruction, we would like to emphasize that they are not the only cause of the bias. Due to a training process driven by photometric loss instead of direct regularization on geometry, it is possible that the reconstructed surfaces still exhibit high reconstruction bias even with fewer LOPs. For example, G-Surfel sometimes shows a high depth error but with fewer LOPs compared to other methods (Fig. 4 (a)).